OpenStudy (anonymous):

simplify i^5 show steps / solution

OpenStudy (anonymous):

the answer is -1 how can you arrive to the answer

myininaya (myininaya):

\[i^5=i^{(4+1)}=i^4*i^1=(1)*i=i\]

myininaya (myininaya):

answer is not -1

OpenStudy (anonymous):

im sorry yeah "i" is the answer

OpenStudy (anonymous):

Yeah, I was gonna say: \[i^2 = -1\]\[\implies i^5 = i^2\cdot i^2 \cdot i = (-1)(-1)i = i\]

OpenStudy (anonymous):

great guys LOL

OpenStudy (anonymous):

\[i=\sqrt{-1} \] \[i^2=\sqrt{-1}\sqrt{-1}=-1\] \[i^4=\sqrt{-1}\sqrt{-1}\sqrt{-1}\sqrt{-1}=(-1)(-1)\] so therefore \[i^5=i\]

myininaya (myininaya):

this is what i always do: say we have \[i^a\] and a=4*b+r \[i^a=i^{4*b+r}=i^{4*b}*i^r=(i^4)^b*i^r=1^b*i^r=1*i^r=i^r\]

OpenStudy (anonymous):

got it tnx

myininaya (myininaya):

if r=0 (if the remainder is 0) then answer is 1 if r=1 (if the remainder is 1) then the answer is i if r=2 (if the remainder is 2) then the answer is -1 if r=3 (if the remainder is 3) then the answer is -i

OpenStudy (anonymous):

^i like

myininaya (myininaya):

number theory is your friend

OpenStudy (anonymous):

The answer is 1 =]

myininaya (myininaya):

i

OpenStudy (anonymous):

lol my b

OpenStudy (anonymous):

i tried to read number theory very think

OpenStudy (anonymous):

thick

myininaya (myininaya):

very think?

myininaya (myininaya):

lol

myininaya (myininaya):

math is a huge subject i'm pretty sure each of the subjects in math are also huge

myininaya (myininaya):

we can't no everything

myininaya (myininaya):

know*

OpenStudy (anonymous):

it was just super in depth,

OpenStudy (anonymous):

nah its all good i am religious but not narrow minded, ive heard these things as well, and also in physics, the movie what the bleep do we know says things like this as well. but i agree with you that i have heard this as well from other sources

OpenStudy (anonymous):

nah i hear you i dont dislike anyone for having different religious views

myininaya (myininaya):

:)

OpenStudy (anonymous):

alright myini have a good night answering peoples questions cause i STRUGGLED TONIghT

myininaya (myininaya):

lol bye bnut